# Analytical approach for calculating shadow of dynamical black hole

### v1

@@ II.2 Dynamical case
$\frac{d}{dv}\left(E^{2}\right)$ $=$ $-\frac{\Lambda}{r},$
[…]
Given that $\Lambda\propto\dot{M}$, the generalized flux is directly related to the rate of change of the black hole’s mass. When the spacetime transitions back to a static configuration, $\dot{M}\to 0$, implying that $\Lambda\to 0$. In this static limit, the dynamics simplify, and the energy per unit mass of the particle becomes conserved. This leads to: $\frac{dE}{d\lambda}\to 0\quad\Rightarrow\quad E^{2}=\text{constant}.$ In this scenario, the geodesic equation reduces to the form seen in a static spacetime, recovering the standard result for conserved energy, as in (II.2).
The total generalized flux $\Lambda$ can indeed be associated with the change in energy, offering insights into the relationship between energy dynamics and energy conditions. Specifically:
  • For a dissipative process, where energy decreases over time ($\dot{E}<0$), we have $\Lambda>0$, implying $\dot{M}>0$. This corresponds to an accretion process, where the black hole gains mass. In this case, the null energy condition (NEC) is satisfied.
  • Conversely, if the energy grows over time ($\dot{E}>0$), then $\Lambda<0$, which implies $\dot{M}<0$. This situation reflects a radiative process where the black hole loses mass. Here, the NEC is violated, as the energy flux is outward.
It is also worth noting that the second-order geodesic equations for timelike or spacelike trajectories differ from those for null trajectories by the inclusion of an additional term: the Newtonian attractive force, $a_{n}$.
The general form of the radial geodesic equation can be written as:
$\frac{d^{2}r}{d\lambda^{2}}=2E\dot{E}-\delta a_{n}+a_{g}+a_{c}.$ (31)
This equation unifies the contributions from energy changes, Newtonian gravity, relativistic corrections, and centrifugal forces, offering a comprehensive description of geodesic motion in both static and dynamical spacetimes.
III Dynamical black hole shadow
III.1 Dynamical photon sphere
To determine the radius of a dynamical photon sphere, we make the following two assumptions:
@@ III.1 Dynamical photon sphere
    Slow Rate of Accretion: The generalized flux is small, $\Lambda\ll 1$, implying that the dynamical changes in the black hole’s mass are slow.
@@
    Radial Acceleration is Zero: The photon sphere is defined by equilibrium, i.e., $\frac{d^{2}r}{d\lambda^{2}}=0$.
To calculate the radius of the photon sphere, we introduce a book-keeping parameter $\alpha$ to facilitate an expansion. This parameter allows us to systematically account for small perturbations. We assume the radius of the photon sphere can be expressed as:
$r_{ph}=r_{0}+\alpha r_{1}(v),$ (32)
where:
  • $r_{0}$ is the radius of the photon sphere in the corresponding static spacetime.
  • $r_{1}(v)$ is a time-dependent perturbation that accounts for the dynamical effects introduced by the time dependence of the spacetime metric.
After performing the necessary expansions and calculations, the parameter $\alpha$ will be set to unity to recover the physical radius. This approach enables a perturbative treatment of the photon sphere’s radius in slowly evolving black hole spacetimes.
The geodesic equation (22) can be written as:
$\alpha a_{i}+a_{g}+a_{c}=0,$ (33)
and expanding with respect to $\alpha$, we find that $r_{0}$ satisfies the algebraic equation:
$a_{g}+a_{c}=0,$ (34)
and the first-order correction is given by:
$r_{1}=-\frac{a_{i}}{a^{\prime}_{g}+a^{\prime}_{c}}.$ (35)
In the paper Vertogradov and Övgün (2024b), we proved that if the weak energy condition is satisfied, then the radius of the photon sphere grows. Let us prove that $r_{1}<0$ only if the weak energy condition is violated. From the definition (21) and $a_{i}$, one can see that the weak energy condition demands $a_{i}\leq 0$ ($\dot{M}\geq 0$). Thus, we need to prove that the denominator of (35) is positive. From the definitions of $a_{g}$ and $a_{c}$, we find:
$a^{\prime}_{g}$ $=$ $\frac{L^{2}}{r^{5}}\left(M^{\prime\prime}r^{2}-6M^{\prime}r+12M\right),$
$a^{\prime}_{c}$ $=$ $-\frac{3L^{2}}{r^{4}}.$ (36)
Thus, one should demand the following inequality  [^1] Remember that we estimate (III.1) at $r=r_{0}$:
$M^{\prime\prime}r^{2}-6M^{\prime}r+12M-3r\geq 0.$ (37)
The dominant energy condition demands $-\frac{M^{\prime\prime}}{r}\leq 2\frac{M^{\prime}}{r^{2}}$. Applying this condition, we obtain:
$-8M^{\prime}r+12M-3r\geq 0$ (38)
The radius of a photon sphere should satisfy the following condition:
$6M-2M^{\prime}r-2r=0,$
$12M-3r-8M^{\prime}r=5r-4M\geq 0$ (39)
The last inequality holds because $r_{0}\geq\frac{3}{2}r_{h}$, where $r_{h}=2M(r_{h},v)$ is the radius of an apparent horizon. Thus, we prove that the denominator of (35) is always positive. This implies that if the weak energy condition is satisfied, then the radius of the photon sphere grows over time, which coincides with the earlier results obtained in Vertogradov and Övgün (2024a).
As an example, let us consider the simplest case of the Vaidya spacetime, which is given by:
$ds^{2}=-\left(1-\frac{2M(v)}{r}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2}.$ (40)
The corresponding static spacetime is the Schwarzschild spacetime, for which the photon sphere radius is $r_{0}=3M_{0}$, where $M_{0}$ is the mass of the Schwarzschild black hole. In the case of accretion ($\dot{M}>0$), the induced acceleration satisfies $a_{i}<0$.
The derivatives of the relativistic correction $a_{g}$ and the centrifugal force $a_{c}$ are given by:
$a^{\prime}_{g}$ $=$ $\frac{12M(v)}{(3M_{0})^{5}}L^{2},$
$a^{\prime}_{c}$ $=$ $-\frac{3L^{2}}{(3M_{0})^{4}}.$ (41)
The first-order correction to the photon sphere radius, $r_{1}$, is then:
$r_{1}=-\frac{(3M_{0})^{5}a_{i}}{12M(v)L^{2}-9M_{0}L^{2}}\geq 0.$ (42)
One should note, that we consider an accretion process onto a black hole of the mass $M_{0}$. That’s why, $M(v)\geq M_{0}$ and the denominator is always positive. Thus, in the case of accretion, the radius of the photon sphere, given by:
$r_{ph}=r_{0}+r_{1}(v),$ (43)
is always increasing shown in Fig. 1. This result demonstrates that accretion leads to an outward shift of the photon sphere in the Vaidya spacetime.
Figure 1: The Figure shows the variation of $r_{ph}$ due to the parameter $M(v)$ with constant $L=1$ and $a_{i}=-0.1$. Hence $r_{ph}$ remains positive throughout, indicating a increase in the black hole’s photon sphere.
III.2 Dynamical shadow
In order to calculate the angular radius of a shadow seen by an observer, we use the tetrad field, which defines the local orthonormal basis for the observer. The components of the tetrad field are given by:
$e_{0}$ $=$ $\frac{1}{\sqrt{1-\frac{2M}{r}}}\frac{d}{dv},$
$e_{1}$ $=$ $e_{0}+\sqrt{1-\frac{2M}{r}}\frac{d}{dr},$
$e_{2}$ $=$ $\frac{1}{r}\frac{d}{d\theta},$
$e_{3}$ $=$ $\frac{1}{r\sin\theta}\frac{d}{d\varphi}.$ (44)
In the equatorial plane ($\theta=\frac{\pi}{2}$), the null geodesic is described by the coordinates $v(\lambda)$, $r(\lambda)$, and $\varphi(\lambda)$. The tangent vector to the null geodesic can be expressed as:
$\dot{v}\frac{d}{dv}+\dot{r}\frac{d}{dr}+\dot{\varphi}\frac{d}{d\varphi}=\xi\left(e_{0}+\cos\omega\,e_{1}-\sin\omega\,e_{3}\right).$ (45)
Here, $\xi$ is a scalar factor, and $\omega$ is the angle between the null geodesic and the radial direction in the observer’s rest frame. By comparing the coefficients of $\frac{d}{dr}$ and $\frac{d}{d\varphi}$, we can write:
$\dot{r}$ $=$ $\sqrt{1-\frac{2M}{r}}\cos\omega,$
$\dot{\varphi}$ $=$ $-\frac{1}{r}\sin\omega.$ (46)
To proceed further, we need an expression for $\frac{dr}{d\varphi}$. Using our ansatz (II.2), along with (III.2), and (14), we find:
$\left(\frac{dr}{d\varphi}\right)^{2}$ $=$ $r^{2}f\cot^{2}\omega,$
$\sin^{2}\omega_{sh}$ $=$ $\frac{L^{2}}{E^{2}(v)}\frac{f\left(r_{o}\right)}{r_{o}^{2}},$ (47)
where $\sin^{2}\omega_{sh}$ represents the angular size of the shadow as seen by the observer located at $r_{o}$.
Hence, the angular size of the shadow decreases if $\dot{E}>0$ and increases if $\dot{E}<0$. One can assert the following: In a dynamical black hole, the shadow grows if light loses energy while propagating near the black hole and decreases if light gains energy while propagating near the black hole. We know, that an observer obtains light if the following condition is held
$\frac{L^{2}}{E^{2}(v)}\geq\frac{r^{2}_{ph}}{f(r_{ph})}.$ (48)
Hence, we can estimate the angular size of a shadow as
$\sin^{2}\omega_{sh}=\frac{r^{2}_{ph}f(r_{o})}{r_{o}^{2}f(r_{ph})}.$ (49)
Also one can write
$E^{2}(v)=E^{2}_{0}+\mu(v).$ (50)
Here $E_{0}$ is initial energy of a particle and $\mu(v)$ is an arbitrary function of time. $\mu(v)>0$ if null energy condition is violated and $\mu(v)<0$ if it is held. Denoting $\gamma(v)\equiv\frac{\mu(v)}{E^{2}_{0}}$, one can write
$\sin^{2}\omega_{sh}=\frac{b_{0}^{2}}{1+\gamma(v)}\frac{f(r_{0})}{r_{0}^{2}}.$ (51)
Here $b_{0}$ is a shadow radius of corresponding static spacetime.
For $r_{o}\gg M$, the term $\frac{2M}{r_{o}}$ becomes negligible compared to $1$. Hence, the metric function at the observer radius simplifies as: $f(r_{o})=1-\frac{2M}{r_{o}}\approx 1$. Substituting this approximation into the expression for $\sin^{2}\omega_{sh}$ gives approximately:
$\omega_{sh}\approx\frac{3}{r_{o}}\sqrt{\frac{3M}{1-\nu v}\cdot}.$ (52)
For Vaidya spacetime one can take $b_{0}^{2}=27M$ and $\gamma(v)=-\nu v,~{}~{}\nu>0$.
The plot in 2 shows that as $v$ increases, $\omega_{sh}$ grows due to the factor $(1-\nu v)^{-1/2}$, reflecting the time-dependence introduced by the Vaidya spacetime.
Figure 2: The Figure shows the plot of $\omega_{sh}(v)$ as a function of the retarded time $v$ for the Vaidya spacetime, using the approximation for $r_{o}\gg M$. The parameters are $r_{o}=10$, $M=1$, and $\nu=0.1$.
Figure 3: The Figure shows the visualization of the black hole shadow for different combinations of mass $(M=10,20)$ and retarded time $(v=2,5)$. Each subplot corresponds to a specific pair of $M$ and $v$, showing how the shadow radius change with these parameters.
The figure 3 illustrates the black hole shadow for varying combinations of mass $(M)$ and retarded time $(v)$, showcasing how these parameters influence the shadow’s size. As $M$ increases, the shadow radius grows significantly, consistent with the theoretical expectation that the shadow size scales with the black hole’s mass due to stronger gravitational lensing. For a fixed $M$, an increase in $v$ slightly enlarges the shadow radius, reflecting the dynamic nature of the Vaidya spacetime, where the metric evolves with time $v$. The mass has a dominant effect on the shadow size, while the time evolution modulates it further, demonstrating the interplay between static and dynamic factors in determining the observable shadow characteristics. This highlights the potential to use shadow observations to probe both the mass and dynamical properties of black holes.
IV Conclusion
A real astrophysical black hole is surrounded by an accretion disc. As a result of constant accretion onto the black hole, its mass changes over time. Consequently, the spacetime around the black hole becomes time-dependent. This time-dependence should, in turn, influence observable manifestations of the black hole, such as its shadow. It has previously been shown that the evolution of the shadow is intricately connected to the energy conditions of the spacetime Vertogradov and Övgün (2024b). Variations in the mass of the black hole, driven by accretion or radiation, directly affect the size and shape of the shadow, making it a potential tool to study dynamic processes in astrophysical black holes.
Recently, it has been shown that the shadow also depends on the structure of the apparent horizon Vertogradov (2024a). This connection implies that by studying the dynamical shadow, we can gain significant insights into the processes occurring near the horizon of a black hole. However, the analysis of dynamical processes is inherently challenging due to the absence of a conserved energy for the particle in a time-dependent spacetime. This lack of conserved energy complicates the equations of motion, preventing the reduction of the second-order equations to first-order ones, thereby making the study of dynamical black holes mathematically more complex.
In our work, we analyzed the second-order geodesic equation and expressed the radial component in terms of fictitious forces. In a static metric, the radius of the photon sphere is determined by the balance between the centrifugal force and the general relativistic corrections. By extending this framework to a dynamical spacetime, we assumed that the radius of the dynamical photon sphere is established by the balance of three forces: the induced force (arising from the black hole’s mass variation), the centrifugal force, and the general relativistic corrections. Using this assumption, we derived an approximate expression for the evolution of the photon sphere radius, capturing its dependency on the dynamic properties of the black hole and the surrounding spacetime.
Then, we examined the formation of the shadow of a dynamical black hole, expressing the change in the energy of light through the generalized flux. We demonstrated that the shadow of the black hole increases if the light loses energy and decreases if the light gains energy during its propagation near the black hole.
An intriguing observation is that when the radius of the photon sphere coincides with the radius of the NEC (Null Energy Condition) horizon Vertogradov (2024a), the dynamical shadow does not evolve, even though both the photon sphere and the NEC horizon radii are changing. This result highlights a surprising decoupling between the evolution of these radii and the observable shadow in this specific configuration.
We have developed a method for calculating the dynamical shadow of a black hole under the assumption of a small accretion rate. From an astrophysical perspective, the changes caused by accreting matter typically do not significantly affect the mass of the black hole. Consequently, the evolution of the photon sphere radius and the shadow of the black hole is expected to be slow and largely undetectable by modern experimental tools.
The proposed method represents a foundational step toward the analytical calculation of the shadow of a dynamical black hole in the general case. Its refinement will enable the description of the shadow in more extreme scenarios, such as the gravitational collapse of a star—a process so rapid that the evolution and formation of the shadow could become observable. This capability would provide profound insights into the dynamic processes involved in black hole formation and their observational signatures.

### v2

@@ II.2 Dynamical case
$\dot{\varphi}$ $=\frac{L}{r^{2}},$ (11)
$-\frac{1}{2}f^{\prime}\,\dot{v}^{2}+r\,\dot{\varphi}^{2}-\varepsilon\,\ddot{v}$ $=0,$ (12)
$\varepsilon\,\ddot{r}=\frac{1}{2}\dot{f}\,\dot{v}^{2}+f\,\ddot{v}+f^{\prime}\,\dot{v}\,\dot{r}.$ (13)
Eq. (12) links curvature gradients ($f^{\prime}$) to the time acceleration $\ddot{v}$, while (13) isolates the radial dynamics. The null condition $f\dot{v}^{2}+2\varepsilon\,\dot{v}\,\dot{r}+L^{2}/r^{2}=0$ and (12) imply
$f\,\ddot{v}=\varepsilon\,f\,\frac{L^{2}}{r^{3}}-\frac{\varepsilon}{2}ff^{\prime}\,\dot{v}^{2},$ (14)
and
$f^{\prime}\,\dot{v}\,\dot{r}=\frac{1}{2\varepsilon}ff^{\prime}\,\dot{v}^{2}-\frac{1}{2\varepsilon}\frac{f^{\prime}L^{2}}{r^{2}}.$ (15)
Taken together they remove $\ddot{v}$ and $\dot{v}\,\dot{r}$ from (13), exposing the physical forces. Substituting (14)-(15) into (13) and using f=12M/rf=1-2M/r yields
$\ddot{r}=-\,\varepsilon\,\frac{\dot{M}}{r}\,\dot{v}^{2}+\Big(M^{\prime}r-3M\Big)\frac{L^{2}}{r^{4}}+\frac{L^{2}}{r^{3}}\equiv a_{i}+a_{g}+a_{c}.$ (16)
The new ingredient $a_{i}=-\varepsilon\,(\dot{M}/r)\dot{v}^{2}$ is an induced acceleration driven by accretion ($\dot{M}>0$) or mass loss ($\dot{M}<0$). It vanishes in the static limit and has the correct sign to oppose (assist) infall for accretion (evaporation/outflow). Note that (i) If $\dot{M}=0$ we recover the static result (6) with $\delta=0$. (ii) For pure Vaidya $M=M(v)$, $M^{\prime}=0$ and the GR term simplifies to $3ML^{2}/r^{4}$. (iii) Reparametrizations $\lambda\to a\lambda+b$ leave (16) homogeneous, as required. Define the canonical energy (the momentum conjugate to $v$)
$E(v,r)\equiv-\frac{\partial\mathcal{L}}{\partial\dot{v}}=f(v,r)\,\dot{v}-\varepsilon\,\dot{r},$ (17)
and the null effective potential identity
$\Big(\frac{dr}{d\lambda}\Big)^{2}+V_{\rm eff}(v,r;E)=0,\qquad V_{\rm eff}=f(v,r)\frac{L^{2}}{r^{2}}-E^{2}.$ (18)
Remark. Unlike in the static case, $E$ depends on both $v$ and $r$; nevertheless, its total evolution is governed by a single scalar flux. Differentiating (18) along the geodesic and inserting (16) yields the gauge-invariant law
$\frac{d}{dv}\Big(E^{2}\Big)=-\,\frac{\Lambda}{r},\qquad\Lambda\equiv\varepsilon\,\dot{M}\,\dot{v}^{2},$ (19)
hence
$E^{2}(v,r)=E_{0}^{2}-\int^{v}\frac{\Lambda(v^{\prime},r)}{r}\,dv^{\prime}.$ (20)
$\Lambda$ is a generalized (apparent) flux. For accretion ($\dot{M}>0$) we have $\Lambda>0$ and $dE/dv<0$: light loses energy while propagating in. For mass loss/outflow the sign reverses.
III Dynamical photon sphere and shadow
III.1 Adiabatic photon sphere
We assume an adiabatic regime where the induced term is subleading:
$|a_{i}|\ll\min\{|a_{c}|,|a_{g}|\}\quad\Longleftrightarrow\quad\epsilon_{\rm ad}\equiv\frac{|\Lambda|\,r}{L^{2}}\ll 1.$ (21)
$\epsilon_{\rm ad}$ compares the instantaneous forcing by $\dot{M}$ to the static curvature/centrifugal scales; it is dimensionless and small for mild accretion rates and moderate $L$. At fixed $v$, define $r_{ph}(v)$ by $\ddot{r}=0$. Introducing
$r_{ph}(v)=r_{0}+\alpha\,r_{1}(v),$ (22)
with bookkeeping parameter $\alpha$ (set to unity at the end), the zeroth-order radius $r_{0}$ solves (8), while the first-order shift is
$r_{1}(v)=-\frac{a_{i}(r_{0};v)}{a_{g}^{\prime}(r_{0})+a_{c}^{\prime}(r_{0})},$ (23)
where
$a_{g}^{\prime}=\frac{L^{2}}{r^{5}}(M^{\prime\prime}r^{2}-6M^{\prime}r+12M),\qquad a_{c}^{\prime}=-\frac{3L^{2}}{r^{4}}.$ (24)
Under accretion ($\dot{M}\geq 0$) one has $a_{i}\leq 0$. If $a_{g}^{\prime}+a_{c}^{\prime}>0$ (a condition compatible with standard energy inequalities; cf. $r_{0}\geq 3M$ in Schwarzschild), then $r_{1}\geq 0$ and the photon sphere expands. For example of Vaidya case: For $M=M(v)$, $r_{0}=3M_{0}$ and
$a_{g}^{\prime}=\frac{12M(v)}{(3M_{0})^{5}}L^{2},\qquad a_{c}^{\prime}=-\frac{3L^{2}}{(3M_{0})^{4}},$ (25)
so
$r_{1}(v)=-\,\frac{(3M_{0})^{5}\,a_{i}}{L^{2}\,[\,12M(v)-9M_{0}\,]}\geq 0\quad(\dot{M}\geq 0),$ (26)
and $r_{ph}(v)=r_{0}+r_{1}(v)$ increases monotonically with $M(v)$ in the adiabatic regime. If $M(v)\to M_{0}$ one recovers $r_{1}\to 0$.
III.2 Dynamical shadow
We adopt an EF-static observer in the ingoing chart ($\varepsilon=+1$) with orthonormal basis
$e_{0}=f^{-1/2}\partial_{v},\quad e_{1}=e_{0}+f^{1/2}\partial_{r},\quad e_{2}=r^{-1}\partial_{\theta},\quad e_{3}=(r\sin\theta)^{-1}\partial_{\varphi}.$ (27)
In the equatorial plane, writing the null tangent as $\xi\,(e_{0}+\cos\omega\,e_{1}-\sin\omega\,e_{3})$ gives
$\dot{r}=\sqrt{f}\,\cos\omega,\qquad\dot{\varphi}=-\frac{1}{r}\sin\omega,$ (28)
and combining with $\dot{\varphi}=L/r^{2}$ and (17)-(18) yields
$\sin^{2}\omega_{sh}=\frac{L^{2}}{E(v,r_{o})^{2}}\,\frac{f(r_{o})}{r_{o}^{2}}=\frac{b_{\rm crit}(v)^{2}}{r_{o}^{2}}\,f(r_{o}),\qquad b_{\rm crit}(v)\equiv\frac{L}{E(v,r_{o})}=\sqrt{\frac{r_{ph}(v)^{2}}{f\big(r_{ph}(v)\big)}}.$ (29)
Equation (29) separates propagation (through $E$) from geometry (through $r_{ph}$). Via (19), accretion ($\dot{M}>0$) reduces $E$ and therefore increases the apparent shadow angle. For a distant observer $r_{o}\gg M$, $f(r_{o})\simeq 1$ so $\omega_{sh}\simeq b_{\rm crit}(v)/r_{o}$; in Vaidya $b_{\rm crit}\simeq 3\sqrt{3}\,M(v)$ at leading adiabatic order. Our instantaneous $r_{ph}(v)$ is the spherically symmetric specialization of a time-dependent photon surface in the sense of Claudel–Virbhadra–Ellis. The present adiabatic expansion makes this notion constructive by delivering the explicit first-order shift (23) and the associated impact parameter in (29).
Example: (Adiabatic Vaidya: steady accretion).
Consider $M(v)=M_{0}+\nu v$ with $\nu>0$ and $\nu v\ll M_{0}$ (adiabatic regime), in the ingoing EF chart ($\varepsilon=+1$). To first order in the induced term, the photon-sphere shift derived in Eq. (23) evaluates—using the circular null relation $\dot{r}=0$ at $r_{0}=3M_{0}$—to the compact form
$r_{ph}(v)\simeq 3M_{0}\;+\;r_{1}(v),\qquad r_{1}(v)=\frac{9\,\nu\,M_{0}^{2}}{M_{0}+4\,\nu v},$ (30)
which is positive and slowly decreasing in $v$ (the $\dot{M}$-driven correction weakens as $M$ grows). At leading adiabatic order the critical impact parameter is
$b_{\rm crit}(v)\simeq 3\sqrt{3}\,M(v)=3\sqrt{3}\big(M_{0}+\nu v\big),$ (31)
so for a distant static observer at $r_{o}\gg M(v)$,
$\omega_{sh}(v)\simeq\frac{3\sqrt{3}}{r_{o}}\,\big(M_{0}+\nu v\big).$ (32)
Accretion ($\dot{M}>0$) grows both $r_{ph}$ and $b_{\rm crit}$; the shadow angle increases linearly in $v$ at this order. Set $M_{0}=1$, $\nu=10^{-3}$, $r_{o}=100$ (all in geometrized units). The table reports $M(v)$, $b_{\rm crit}(v)\simeq 3\sqrt{3}M(v)$, and $\omega_{sh}=b_{\rm crit}/r_{o}$:
Table 1: Adiabatic Vaidya with steady accretion. Parameters: $M_{0}=1$, $\nu=10^{-3}$, $r_{o}=100$ (geometrized units). We use $M(v)=M_{0}+\nu v$, $b_{\rm crit}(v)\simeq 3\sqrt{3}\,M(v)$, and $\omega_{sh}\simeq b_{\rm crit}(v)/r_{o}$ (cf. Eq. (32)). The data illustrate the linear growth of the shadow angle with $v$.
$v$ $M(v)$ $b_{\rm crit}(v)$ $\omega_{sh}$ (rad) $\omega_{sh}$ (deg)
0 $1.000$ $5.196$ $0.05196$ $2.978$
$50$ $1.050$ $5.455$ $0.05455$ $3.126$
$100$ $1.100$ $5.716$ $0.05716$ $3.275$
Table 1 shows that Each 5% increase in $M$ produces a $\sim$5% increase in $b_{\rm crit}$ and hence a proportional rise in $\omega_{sh}$, as expected from Eq. (32). The step $\Delta v=50$ keeps $\nu v/M_{0}\ll 1$, ensuring $\epsilon_{\rm ad}\ll 1$ and validating the linearized treatment. For a different observer radius $r_{o}$, multiply the angles by $100/r_{o}$ (since $\omega_{sh}\propto r_{o}^{-1}$). The fractional growth rate is $d\omega_{sh}/dv\simeq(3\sqrt{3}/r_{o})\,\nu$; with the numbers above this is $\simeq 5.196\times 10^{-5}$ rad per unit $v$.
Example 2 (Evaporation/outflow: shrinking shadow).
Keep the ingoing EF chart ($\varepsilon=+1$) but take a slowly decreasing mass $M(v)=M_{0}-\nu v$ with $\nu>0$ and $0\leq v\ll M_{0}/\nu$. Then
$r_{ph}(v)\simeq 3M_{0}\;+\;r_{1}(v),\qquad r_{1}(v)=-\,\frac{9\,\nu\,M_{0}^{2}}{\,M_{0}-4\,\nu v\,}<0,$ (33)
so the photon sphere contracts. At leading order,
$b_{\rm crit}(v)\simeq 3\sqrt{3}\,M(v)=3\sqrt{3}\big(M_{0}-\nu v\big),\qquad\omega_{sh}(v)\simeq\frac{3\sqrt{3}}{r_{o}}\,\big(M_{0}-\nu v\big),$ (34)
and the shadow angle decreases linearly. With $M_{0}=1$, $\nu=10^{-3}$, $r_{o}=100$:
Table 2: Adiabatic Vaidya with slow mass loss. Parameters: $M_{0}=1$, $\nu=10^{-3}$, $r_{o}=100$. Here $M(v)=M_{0}-\nu v$, $b_{\rm crit}(v)\simeq 3\sqrt{3}\,M(v)$, and $\omega_{sh}\simeq b_{\rm crit}(v)/r_{o}$. The entries show the linear decrease of the shadow angle predicted by Eq. (33).
$v$ $M(v)$ $b_{\rm crit}(v)$ $\omega_{sh}$ (rad) $\omega_{sh}$ (deg)
0 $1.000$ $5.196$ $0.05196$ $2.978$
$50$ $0.950$ $4.936$ $0.04936$ $2.828$
$100$ $0.900$ $4.677$ $0.04677$ $2.679$
Table 2 shows that decreasing $M$ reduces $b_{\rm crit}$ and thus $\omega_{sh}$ linearly in $v$, mirroring the accretion case with the opposite sign. Ensure $M(v)>0$ and $\nu v\ll M_{0}$ to remain within the adiabatic regime used to derive Eqs. (33)-(29). For fixed $r_{o}$, relative changes in $\omega_{sh}$ directly trace relative changes in $M$ at leading order. Require $M(v)>0$ and $\epsilon_{\rm ad}=\frac{|\Lambda|\,r}{L^{2}}\ll 1$; numerically this means small $\nu$ and moderate $L$.
As predicted by Eq. (32), the shadow angle grows linearly in $v$ during accretion; see Table 1. Conversely, for a slowly decreasing mass history, Table 2 illustrates a linear shrinkage of the shadow, consistent with Eq. (33).
IV Conclusion
We have presented a unified and minimal description of null geodesics and black-hole shadows in spherically symmetric spacetimes with mild time dependence. The analysis proceeds from the EF action to a clean “force” decomposition of the radial equation,
@@ IV Conclusion
    The dynamical geodesic equation (16) separates induced ($a_{i}$), centrifugal ($a_{c}$), and general-relativistic ($a_{g}$) contributions. The induced term $a_{i}=-\varepsilon(\dot{M}/r)\dot{v}^{2}$ encodes accretion or outflow and vanishes in the static limit.
@@
    A gauge-invariant energy law, Eq. (19), shows that the squared canonical energy evolves according to the generalized flux $\Lambda=\varepsilon\dot{M}\dot{v}^{2}$. Accretion ($\dot{M}>0$) decreases $E$ along the ray; outflow increases it.
  3. 3.
    In the adiabatic regime we obtained a closed first-order expression for the photon-sphere shift, Eq. (23). Under standard conditions the denominator $a_{g}^{\prime}+a_{c}^{\prime}$ is positive, so accretion ($a_{i}\leq 0$) yields $r_{1}\geq 0$ and the photon sphere expands.
  4. 4.
    The observable shadow follows from Eq. (29), which ties the angular radius to the evolving critical impact parameter $b_{\rm crit}(v)=\sqrt{r_{ph}^{2}/f(r_{ph})}$. For distant observers $\omega_{sh}\simeq b_{\rm crit}(v)/r_{o}$. In Vaidya, $b_{\rm crit}\simeq 3\sqrt{3}\,M(v)$ at leading order, providing a simple scaling of the shadow with the mass history.
The framework makes quantitative the intuitive statement that inflow (outflow) enlarges (shrinks) both the photon sphere and the shadow. Because our expressions are algebraic in $M(v,r)$ and its derivatives, they can be combined with parametric or semi-analytic accretion histories to produce observer-ready predictions for time-variable shadow sizes. The treatment assumes spherical symmetry and an adiabatic hierarchy $|a_{i}|\ll\min\{|a_{c}|,|a_{g}|\}$; it neglects spin, plasma dispersion, and radiative transfer through an emitting medium. Within these limits, the results provide controlled leading-order estimates. Immediate extensions include (i) non-adiabatic evolution during rapid collapse, (ii) incorporation of refractive/plasma effects, (iii) generalization to slowly rotating and radiating backgrounds (e.g. Kerr-Vaidya-type metrics), and (iv) comparison with numerical ray tracing and horizon-scale imaging of time-variable sources. The compact relations (23) and (29) are designed to serve as analytic benchmarks for such studies.
Our analysis refines and extends classic force-balance ideas for photon motion Abramowicz et al. (1988); Abramowicz and Prasanna (1990) to dynamical settings, and it complements photon-surface approaches in general spacetimes Claudel et al. (2001). While we focused on adiabatic evolution, the formalism provides a baseline for treating faster processes (e.g. late-stage collapse) and for confronting shadow variability with models of time-dependent inflow or outflow. Future work will relax the adiabatic assumption, incorporate plasma effects, and compare with numerical ray-tracing in fully dynamical geometries.